OpenStudy (anonymous):

1. AD= 12 and CD=8 find BD ***************metal****************

OpenStudy (anonymous):

weres the diagram

OpenStudy (anonymous):

hold up girl ill show it in a bit problem 36

OpenStudy (anonymous):

diagrem next to it

OpenStudy (anonymous):

@jim_thompson5910 help me please :)

OpenStudy (anonymous):

u have to get the altitude ....

jimthompson5910 (jim_thompson5910):

let x = length of BD AD/BD = BD/CD 12/x = x/8 12*8 = x*x 96 = x^2 ... ... ... x = ??

OpenStudy (anonymous):

??\\

jimthompson5910 (jim_thompson5910):

if x^2 = 96, then what is x

OpenStudy (anonymous):

x-48

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

i mean x=48

jimthompson5910 (jim_thompson5910):

you don't divide both sides by 2

jimthompson5910 (jim_thompson5910):

you take the square root of both sides (to undo the square) and you simplify as much as possible \[\large x^2 = 96\] \[\large x = \sqrt{96}\] \[\large x = \sqrt{16*6}\] \[\large x = \sqrt{16}*\sqrt{6}\] \[\large x = 4\sqrt{6}\]

OpenStudy (anonymous):

thnxxx'

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

@jim_thompson5910 can you check the diagram and help me on 38 please

jimthompson5910 (jim_thompson5910):

same idea, but now AD/BD = BD/CD sqrt(2)/x = x/sqrt(2) sqrt(2)*sqrt(2) = x*x (sqrt(2))^2 = x^2 x = ???

OpenStudy (anonymous):

????

jimthompson5910 (jim_thompson5910):

if x^2 = (sqrt(2))^2 then what is x

jimthompson5910 (jim_thompson5910):

x = ??? means I want you to fill in the blank

OpenStudy (anonymous):

sorrry but i dont get it at all

jimthompson5910 (jim_thompson5910):

a square root undoes a square, so squaring a square root makes it go away

jimthompson5910 (jim_thompson5910):

this is why (sqrt(2))^2 = 2 so x^2 = 2 x = sqrt(2)

OpenStudy (anonymous):

by geometric mean property \[BD^2=AD \times CD=12\times 8\] \[BD^2=96 \rightarrow BD = \sqrt{96}=4\sqrt{6}\]

OpenStudy (anonymous):

thnxxxxxx all of u